OpenStudy (anonymous):

xdx+ydy/xdx-ydy=(a^2-x^2-y^2)^1/2/(x^2+y^2)^1/2 solve the differentions eqution

OpenStudy (anonymous):

Can you use the equation button below

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

\[x^2 + y^2 = r\] \[x = rcos\theta\] \[y=rsin\theta\]

OpenStudy (anonymous):

\[tan \theta = \frac{y}{x}\]

OpenStudy (anonymous):

Now\[xdx + ydy=rdr\]

OpenStudy (anonymous):

\[d(\frac{y}{x}) = d(tan\theta)\]

OpenStudy (anonymous):

I am sorry I can't get it

OpenStudy (anonymous):

I can't get denominator in the equation if it had been xdy - ydx then I would have done that